### v1

# The Gödel Universe as a Lie group with left-invariant Lorentz metric and the Iwasawa decomposition

Abstract.
We discuss models of the Gödel Universe as Lie groups with left-invariant Lorentz metric for two simply connected four-dimensional Lie groups, the Iwasawa decomposition for semisimple Lie groups, and left-invariant Lorentz metric on $\operatorname{SL}(2,\mathbb{R})$, following K.-H. Neeb. Also we show that the isometry between two non-isomorphic sub-Riemannian Lie group, constructed by A. Agrachev and D. Barilari, is induced by some Iwasawa decomposition of $\operatorname{SL}(2,\mathbb{R})$.
Keywords and phrases: Gödel Universe, Iwasawa decomposition, left-invariant Lorentz metric, left-invariant sub-Riemannian metric, Lie algebra, Lie group.
The work was carried out within the framework of the State Contract to the IM SB RAS, project FWNF-2022-0006.
1. Introduction
Kurt Gödel in paper [1] of 1949 introduced the Lorentz metric (1) of the signature (+,,,)(+,-,-,-)( + , - , - , - ) on the space $\mathbb{R}^{4}.$ The Gödel Universe (space-time) $S$ is a solution of the General Relativity Theory (the Einstein gravitation equations).
In paper [2], the author found timelike and isotropic geodesics of the Gödel Universe, considered as the Lie group $G=(\mathbb{R},+)\times\operatorname{A}^{+}(\mathbb{R})\times(\mathbb{R},+)$ with left invariant Lorentz metric. Here $\operatorname{A}^{+}(\mathbb{R})$ is connected Lie group of affine transformations on $\mathbb{R}.$
Professor Karl–Hermann Neeb wrote to the author that it is possible to realize the Gödel Universe otherwise and sent an electronic version of his joint with Joachim Hilgert book [3], where in section 2.7 ’’Gödel’s cosmological model and universal covering of $\operatorname{SL}(2,\mathbb{R})$’’ is suggested a left-invariant Lorentz metric on this Lie group.
In this connection, it is useful to mention the paper [4] by A. Agrachev and D. Barilari, where the autors obtained a full classification of left-invariant sub-Riemannian metrics on three-dimensional Lie groups and ’’explicitly find a sub-Riemannian isometry between nonisomorphic Lie groups $\operatorname{SL}(2,\mathbb{R})$ and $\operatorname{SO}(2)\times A^{+}(\mathbb{R})$’’ [4].
The existence of such isometry was indicated ealier in [5] by Falbel and Gorodski.
In a message to the author, Professor Neeb explains this by a diffeomormism of Lie groups $\operatorname{SL}(2,\mathbb{R})$ and $\operatorname{SO}(2)\times A^{+}(\mathbb{R})$ by means of the Iwasawa decomposition for $\operatorname{SL}(2,\mathbb{R})$. Let us cite now only Theorems 6.5.1 and 9.1.3 from [6].
In p. 10.6.4 (i) from [6] are indicated the following isomorphisms of Lie algebras:
$\mathfrak{sl}(2,\mathbb{R})\cong\mathfrak{su}(1,1)\cong\mathfrak{so}(2,1)\cong\mathfrak{sp}(1,\mathbb{R}).$
Consequently, simply connected Lie groups with these Lie algebras are isomorphic.
In Theorem 3 we prove some properties of special left-invariant Lorentz metrics on three-dimensional Lie groups. Also we show in Proposition 3 that the isometry between two non-isomorphic sub-Riemannian Lie group, constructed by A. Agrachev and D. Barilari, is induced by some Iwasawa decomposition of $\operatorname{SL}(2,\mathbb{R})$.
The author expresses his gratitudes to Professor Neeb for fruitful discussions.
2. The Gödel Universe as a Lie group with left-invariant Lorentz metric
Gödel introduced in [1] his space-time $S$ as $\mathbb{R}^{4}$ with the linear element
$ds^{2}=a^{2}\left(dx_{0}^{2}+2e^{x_{1}}dx_{0}dx_{2}+\frac{e^{2x_{1}}}{2}dx_{2}^{2}-dx_{1}^{2}-dx_{3}^{2}\right),\quad a>0.Ыд$ (1)
Gödel noticed in [1] that it is possible to rewrite this quadratic form in view of
$ds^{2}=a^{2}\left[\left(dx_{0}+e^{x_{1}}dx_{2}\right)^{2}-dx^{2}_{1}-\frac{e^{2x_{1}}}{2}dx^{2}_{2}-dx^{2}_{3}\right],$ (2)
which shows obvious that its signature is equal everywhere to (+,,,).(+,-,-,-).( + , - , - , - ) .
We shall assume that $a=1.$
Gödel’s noticed in [1] that on $(S,ds^{2})$ acts simply transitively a four-dimensional isometry Lie group. It is easy to see that such action could be written as
$x_{0}=x_{0}^{\prime}+a,\quad x_{1}=x_{1}^{\prime}+b,\quad x_{2}=x_{2}^{\prime}e^{-b}+c,\quad x_{3}=x_{3}^{\prime}+d$ (3)
with arbitrary $a,b,c,d\in\mathbb{R}.$ This implies that corresponding Lie group $G$ is the simplest simply connected noncommutative four-dimensional Lie group of the view
$G\cong[(\mathbb{R},+)\times G_{2}]\times(\mathbb{R},+),$ (4)
where $G_{2}$ is unique up to isomorphism, necessary isomorphic to $\mathbb{R}^{2}$, two-dimensional noncommutative Lie group. The Lie group $G_{2}$ is isomorphic to the Lie group $A^{+}(\mathbb{R})$ of preserving orientation affine transformations of the real direct line $(\mathbb{R},+).$
[…]
$A=\left(\begin{array}[]{ccccc}e^{-b}&0&0&0&c\\ 0&1&0&0&b\\ 0&0&1&0&a\\ 0&0&0&1&d\\ 0&0&0&0&1\end{array}\right).$ (5)
Under this the equality
$\left(\begin{array}[]{ccccc}e^{-x_{1}}&0&0&0&x_{2}\\ 0&1&0&0&x_{1}\\ 0&0&1&0&x_{0}\\ 0&0&0&1&x_{3}\\ 0&0&0&0&1\end{array}\right)(0,0,0,0,1)^{T}=(x_{2},x_{1},x_{0},x_{3},1)^{T}$ (6)
sets the bijection of the group $G$ onto $\mathbb{R}^{4}$ and the unit of $G$ corresponds to the zero-vector $(0,0,0,0)\in\mathbb{R}^{4}.$ On base of this, (4) and (1), we can identify $(S,ds^{2})$ with the Lie group $G$ equipped with left-invariant Lorentz metric. Let
$e_{0}=\frac{\partial}{\partial x_{0}}(0),e_{1}=\frac{\partial}{\partial x_{1}}(0),e_{2}=\frac{\partial}{\partial x_{2}}(0),e_{3}=\frac{\partial}{\partial x_{3}}(0))$
be the basis of the Lie algebra $\mathfrak{g}$ of the Lie group $G$ at the unit of $G$, corresponding to coordinates $(x_{0},x_{1},x_{2},x_{3})$. Then, according to what has been said and (1), the components of the linear element $ds^{2}$ with respect to this basis are equal to
$g_{00}=1,g_{02}=g_{20}=1,g_{22}=\frac{1}{2},g_{11}=-1,g_{33}=-1,g_{ij}=g_{ji}=0,i\neq j,j\neq 2.$ (7)
According to (6), the Lie subgroup $G_{3}$ can be identified with the matrix Lie group
$\left(\begin{array}[]{cccc}e^{-x_{1}}&0&0&x_{2}\\ 0&1&0&x_{1}\\ 0&0&1&x_{0}\\ 0&0&0&1\end{array}\right),\quad(x_{0},x_{1},x_{2})\in\mathbb{R}^{3}.$ (8)
It is obvious that $(S,ds^{2})=(S_{0},ds_{0}^{2})\times(S_{1},ds_{1}^{2}),$ where $S_{0}=\mathbb{R}^{3},$ $S_{1}=\mathbb{R},$
$ds_{0}^{2}=dx_{0}^{2}+2e^{x_{1}}dx_{0}dx_{2}+\frac{e^{2x_{1}}}{2}dx_{2}^{2}-dx_{1}^{2},\quad ds_{1}^{2}=-dx_{3}^{2}.$ (9)
Also it is clear that we can consider $(S_{0},ds_{0}^{2})$ as the matrix Lie group (8) with left-invariant Lorentz metric, which according to (7) has components
$g_{00}=1,g_{02}=g_{20}=1,g_{22}=\frac{1}{2},g_{11}=-1,g_{ij}=g_{ji}=0,i\neq j,j\neq 2;$ (10)
with respect to the basis $e_{0},e_{1},e_{2}$ of the Lie algebra $\mathfrak{g}_{3}$ of the matrix Lie group (8).
In consequence of (6), for the Lie algebra $\mathfrak{g}_{3}$ of the Lie group $G_{3},$
$e_{0}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\end{array}\right),\quad e_{1}=\left(\begin{array}[]{cccc}-1&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\end{array}\right),\quad e_{2}=\left(\begin{array}[]{cccc}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right).$ (11)
Then in the Lie algebra $\mathfrak{g}_{3}$,
$[e_{1},e_{2}]=e_{1}e_{2}-e_{2}e_{1}=-e_{2},\quad[e_{0},e_{1}]=[e_{0},e_{2}]=0.$ (12)
3. The Iwasawa decompositions of Lie algebras and Lie groups
Let $\mathfrak{g}$ be a semisimple real Lie algebra, $\sigma$ be some Cartan involution of $\mathfrak{g},$ and $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be the corresponding Cartan decomposition ($\mathfrak{k}$ is the Lie subalgebra of $\mathfrak{g}$, consisting of fixed points relative to $\sigma$). Let us denote by $\mathfrak{a}$ a maximal commutative subspace in $\mathfrak{p}.$ Then there is the following Iwasawa decomposition of Lie algebra $\mathfrak{g}$.
Theorem 1.
(4.7.2) in [7]. Let $\mathfrak{g}$ be a semisimple real Lie algebra.
Then there exists a direct sum of vector subspaces in $\mathfrak{g}$
$\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{a}\oplus\mathfrak{n},$ (13)
where $\mathfrak{n}$ is a nilpotent subalgebra in $\mathfrak{g}$ such that the endomorphism $\operatorname{ad}X$ is nilpotent for every $X\in\mathfrak{n},$ and $\mathfrak{a}\oplus\mathfrak{n}$ is a solvable subalgebra in $\mathfrak{g}.$
As an example, the authors of [7] give the decomposition (13) for $\mathfrak{g}=\mathfrak{sl}(n,\mathbb{R}).$ In this case $\mathfrak{k}$ is the Lie subalgebra of skew-symmetric matrices, $\mathfrak{a}$ is the Lie subalgebra of diagonal matrices (with zero trace), and $\mathfrak{n}$ is the Lie subalgebra of strictly upper triangular matrices. In particular, for $\mathfrak{g}=\mathfrak{sl}(2,\mathbb{R})$ we have
$\mathfrak{k}=\left\{\left(\begin{array}[]{cc}0&t\\ -t&0\end{array}\right)\right\},\quad\mathfrak{a}=\left\{\left(\begin{array}[]{cc}t&0\\ 0&-t\end{array}\right)\right\},\quad\mathfrak{n}=\left\{\left(\begin{array}[]{cc}0&t\\ 0&0\end{array}\right)\right\},\quad t\in\mathbb{R},$ (14)
with natural basis
$f_{0}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),\quad f_{1}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),\quad f_{2}=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)$ (15)
and Lie brackets for this basis
$[f_{0},f_{1}]=2f_{0}-3f_{2},\quad[f_{0},f_{2}]=f_{1},\quad[f_{1},f_{2}]=2f_{2}.$ (16)
Let $K=\exp(\mathfrak{k}),$ $A=\exp(\mathfrak{a}),$ $N=\exp(\mathfrak{n})$ be Lie subgroups of the semisimple Lie group $G,$ corresponding to the decomposition (13).
Theorem 2.
(Theorem 9.1.3 in [6]) Let $G$ be a connected semisimple real Lie group. Then $G=KAN$ and the mapping
$(k,a,n)\rightarrow kan$ (17)
is the diffeomorphism of manifold $K\times A\times N$ onto the Lie group $G.$
Corollary 1.
The Lie group $G$ is diffeomorphic to Lie groups $K\times AN$ and $K\times A\times N.$
Theorem 1 and Theorem 6.1.1 from [6] imply the following
Proposition 1.
The sets $K,$ $A,$ $N,$ and $AN$ are connected closed Lie subgroups of the Lie group $G,$ where $\operatorname{Ad}_{G}(K)$ is compact, $A$ is commutative, $N$ is nilpotent, and $AN$ is solvable. The subgroup $K$ contains the center $Z$ of the Lie group $G.$ In addition, $K$ is compact if and only if the center $Z$ of $G$ is finite; in this case $K$ is a maximal compact subgroup of the Lie group $G.$
The statements above imply the following Corollary.
Corollary 2.
If $G=\operatorname{SL}(n,\mathbb{R}),$ then $K=\operatorname{SO}(n),$ $A$ is the group of all real diagonal $(n\times n)$-matrices with unit determinant, $N$ could be considered as the group of all real upper triangular $(n\times n)$-matrices with units on the main diagonal, and $\operatorname{Sol}(n):=AN$ as the group of all real upper triangular $(n\times n)$-matrices with unit determinant.
Corollary 3.
The Lie group $\operatorname{SL}(n,\mathbb{R})$ is diffeomorphic to Lie groups $\operatorname{SO}(n)\times AN$ and $\operatorname{SO}(n)\times A\times N.$ As a consequence, $\operatorname{SL}(2,\mathbb{R})$ is diffeomorphic to the commutative Lie group $\operatorname{SO}(2)\times A\times N.$
4. Applications of Iwasawa decompositions to $\mathfrak{sl}(2,\mathbb{R})$ and $\operatorname{SL}(2,\mathbb{R})$
This is a preparatory section.
Proposition 2.
There exist an isomorphism of the Lie group $G_{3}$ onto the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ and corresponding realization of $(S_{0},ds^{2}_{0})$ as Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric.
Proof.
Comparing (12) and (16), we see that the linear map $\varphi:\mathfrak{a}\oplus\mathfrak{n}\rightarrow\mathfrak{g}$ such that
$\varphi\left(\frac{-f_{1}}{2}\right)=e_{1},\quad\varphi(f_{2})=e_{2}$ (18)
is an isomorphism of Lie algebras. Then (18) defines isomorphism of Lie groups $\psi:\operatorname{Sol}(2)\rightarrow G_{2}:$
$\psi\left(\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)\right)=\left(\begin{array}[]{cccc}e^{-s}&0&0&e^{-s}r\\ 0&1&0&s\\ 0&0&1&0\\ 0&0&0&1\end{array}\right),$ (19)
$\psi\left(\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\right)=\left(\begin{array}[]{cccc}e^{-s}&0&0&r\\ 0&1&0&s\\ 0&0&1&0\\ 0&0&0&1\end{array}\right).$ (20)
We can consider $(\mathbb{R},+)$ as a Lie algebra and as a Lie group. Then the mappings
$t\in(\mathbb{R},+)\rightarrow\left(\begin{array}[]{cc}0&t\\ -t&0\end{array}\right),\quad t\in(\mathbb{R},+)\rightarrow\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right)$ (21)
are correspondingly the isomorphism of Lie algebras and respective universal covering epimorpism of Lie groups. Then there exists unique isomorphism $\psi$ of the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ onto the Lie group $G_{3},$ with properties (19), (20), and $\psi(t)=t$ for $t\in(\mathbb{R},+).$ It follows from previous considerations that we shall realize $(S_{0},ds^{2}_{0})$ as the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric if components of this metric in the basis $\{f_{0},-f_{1}/2,f_{2}\}$ of its Lie algebra will be as in (10).
The corresponding orthonormal basis is
$X=f_{0},\quad Y=-f_{1}/2,\quad Z=\sqrt{2}(f_{0}-f_{2}).$ (22)
Now we shall consider the Iwasawa decompositions for Lie groups $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ and $\operatorname{SL}(2,\mathbb{R})$ and corresponding mutual expressions of partial derivatives of these Lie groups by parameters $r,s,t$ and natural left-invariant vector fields on the above Lie groups which represent some interest.
Let $kan=kan(t,s,r)$ denotes the following matrix from $\operatorname{SL}(2,\mathbb{R}):$
$\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right)\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)=$
$\left(\begin{array}[]{cc}\cos te^{-s/2}&r\cos te^{-s/2}+\sin te^{s/2}\\ -\sin te^{-s/2}&-r\sin te^{-s/2}+\cos te^{s/2}\end{array}\right).$
Then we have the following partial derivatives
$(kan)^{\prime}_{t}=\left(\begin{array}[]{cc}-\sin te^{-s/2}&-r\sin te^{-s/2}+\cos te^{s/2}\\ -\cos te^{-s/2}&-r\cos te^{-s/2}-\sin te^{s/2}\end{array}\right),$ (23)
$(kan)^{\prime}_{s}=\frac{-1}{2}\left(\begin{array}[]{cc}\cos te^{-s/2}&r\cos te^{-s/2}-\sin te^{s/2}\\ -\sin te^{-s/2}&-r\sin te^{-s/2}-\cos te^{s/2}\end{array}\right),$ (24)
$(kan)^{\prime}_{r}=\left(\begin{array}[]{cc}0&\cos te^{-s/2}\\ 0&-\sin te^{-s/2}\end{array}\right).$ (25)
The left-invariant vector fields on $\operatorname{SL}(2,\mathbb{R}),$ corresponding to $f_{0},$ $-f_{1}/2,$ and $f_{2},$ are
$v_{0}=(kan)\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)=\left(\begin{array}[]{cc}-(r\cos te^{-s/2}+\sin te^{s/2})&\cos te^{-s/2}\\ -(-r\sin te^{-s/2}+\cos te^{s/2})&-\sin te^{-s/2}\end{array}\right),$ (26)
$v_{1}=\frac{1}{2}(kan)\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}-\cos te^{-s/2}&r\cos te^{-s/2}+\sin te^{s/2}\\ \sin te^{-s/2}&-r\sin te^{-s/2}+\cos te^{s/2}\end{array}\right),$ (27)
$v_{2}=(kan)\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)=\left(\begin{array}[]{cc}0&\cos te^{-s/2}\\ 0&-\sin te^{-s/2}\end{array}\right)=(kan)^{\prime}_{r}.$ (28)
Comparing (23), (24), (25) and (26), (27), (28), we see that
$(kan)^{\prime}s=v_{1}-rv_{2},\quad(kan)^{\prime}_{t}=e^{-s}(v_{0}-2rv_{1}+(e^{2s}+r^{2}-1)v_{2}).$ (29)
The equalities imply that
$v_{2}=(kan)^{\prime}_{r},v_{1}=(kan)^{\prime}_{s}+r(kan)^{\prime}_{r},v_{0}=e^{s}(kan)^{\prime}_{t}+2r(kan)^{\prime}_{s}+(r^{2}+1-e^{2s})(kan)^{\prime}_{r}$ (30)
Let us consider now the Lie group $K\times AN=\operatorname{SO}(2)\times\operatorname{Sol}(2)$ with the Lie algebra $\mathfrak{k}\oplus(\mathfrak{a}+\mathfrak{n})$ and the same basis $(f_{0},-f_{1}/2,f_{2}).$ Then
$k(t)=\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right),\quad an(s,r)=\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)=\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right),$
$k^{\prime}(t)=\left(\begin{array}[]{cc}-\sin t&\cos t\\ -\cos t&-\sin t\end{array}\right),(an)^{\prime}_{s}=\frac{-1}{2}\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&-e^{s/2}\end{array}\right),(an)^{\prime}_{r}=\left(\begin{array}[]{cc}0&e^{-s/2}\\ 0&0\end{array}\right),$
$v_{0}=\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right)\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)=\left(\begin{array}[]{cc}-\sin t&\cos t\\ -\cos t&-\sin t\end{array}\right)=k^{\prime}(t),$
$v_{1}=\frac{1}{2}\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{cc}-e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right),$
$v_{2}=\left(\begin{array}[]{cc}e^{-s/2}&re^{-s/2}\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)=\left(\begin{array}[]{cc}0&e^{-s/2}\\ 0&0\end{array}\right)=(an)^{\prime}_{r},\quad(an)^{\prime}s=v_{1}-rv_{2},$
$k^{\prime}(t)=v_{0},\quad(an)^{\prime}s=v_{1}-rv_{2},\quad(an)^{\prime}_{r}=v_{2},$ (31)
$v_{0}=k^{\prime}(t),\quad v_{1}=(an)^{\prime}s+r(an)^{\prime}_{r},\quad v_{2}=(an)^{\prime}_{r}.$ (32)
5. Left-invariant Lorentz metrics on $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ and $\operatorname{SL}(2,\mathbb{R})$
Theorem 3.
the Lorentz metric on $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ from [3] is not isometric to the subspace $(S_{0},ds_{0}^{2})$ of the Gödel Universe and the Iwasawa diffeomorphism of $\operatorname{SL}(2,\mathbb{R})$ onto $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ is not isometry for Lorentz metric on $\operatorname{SL}(2,\mathbb{R})$ from [3].
Proof.
For any (pseudo)-Riemannian manifold $M$ with (pseudo)-metric tensor (,),($\cdot$,$\cdot$),( ⋅ , ⋅ ) , the Levi-Civita connection $\nabla$, and smooth vector fields $X,$ $Y,$ $Z$ is valid the following equation (3.5.(7)) from [8]:
$(\nabla_{X}Y,Z)=\frac{1}{2}[X(Y,Z)+Y(Z,X)-Z(X,Y)+(Z,[X,Y])+(Y,[Z,X])-(X,[Y,Z])].$ (33)
As a consequence, if $(M,(\cdot,\cdot))$ is a Lie group $G$ with left-invariant (pseudo)-metric (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) and $X,$ $Y,$ $Z$ are left-invariant, then
$(\nabla_{X}Y,Z)=\frac{1}{2}[(Z,[X,Y])+(Y,[Z,X])-(X,[Y,Z])].$ (34)
Let assume now that $G=\operatorname{SO}(2)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric and $(X,Y,Z)$ be an orthonormal basis of left-invariant vector fields such that $X$ be tangent to $\operatorname{SO}(2)$ and $(X,X)=1.$ Then $Y=\alpha X+Y_{1},$ $Z=\beta X+Z_{1},$ where $Y_{1},Z_{1}\in\mathfrak{sol}(2)$ and
$(\nabla_{X}Y,Z)=\frac{-1}{2}(X,[Y,Z])=\frac{-1}{2}(X,W_{1}):=-\gamma_{1},\quad W_{1}\in\mathfrak{n},$ (35)
$(\nabla_{X}Y,X)=0,\quad(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=\gamma_{1}Z.$
We want to calculate
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y.$
It follows from (34) that
$(\nabla_{Y}Y,X)=(\nabla_{Y}Y,Y)=0,\quad(\nabla_{Y}Y,Z)=-(Y,W_{1}):=-\gamma_{2},\quad\nabla_{Y}Y=\gamma_{2}Z,$
$(\nabla_{X}Z,X)=(\nabla_{X}Z,Z)=0,\quad(\nabla_{X}Z,Y)=\frac{-1}{2}(X,[Z,Y])=\gamma_{1},\quad\nabla_{X}Z=-\gamma_{1}Y,$
$\quad(\nabla_{Y}Z,X)=\gamma_{1},\quad(\nabla_{Y}Z,Y)=\gamma_{2},\quad(\nabla_{Y}Z,Z)=0,\quad\nabla_{Y}Z=\gamma_{1}X-\gamma_{2}Y,$
$R(X,Y)Y=\nabla_{X}(\gamma_{2}Z)-\nabla_{Y}(\gamma_{1}Z)=-\gamma_{1}\gamma_{2}Y-\gamma_{1}^{2}X+\gamma_{1}\gamma_{2}Y=-\gamma_{1}^{2}X.$
Let us calculate yet an analogue of the sectional curvature $(R(X,Y)Y,X)=-\gamma_{1}^{2}.$
In [3], the authors consider the Lie group $\operatorname{SL}(2,\mathbb{R})$ with left-invariant Lorentz metric and orthonormal basis of opposite to above signature of the form
$X=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),Y=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),Z=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\in\mathfrak{sl}(2,\mathbb{R}).$ (36)
Then $[Y,Z]=2\sqrt{2}X$ in $\mathfrak{sl}(2,\mathbb{R}),$ but $[Y,Z]=0$ in the Lie algebra $\mathfrak{so}(2)\oplus\mathfrak{sol}(2),$ and according to the previous calcultions, $\gamma_{1}=0$ in $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ and
$R(X,Y)Y=0.$ (37)
Now let us calculate $R(X,Y)Y$ and $(R(X,Y)Y,X)$ for left-invariant Lorentz metric on the Lie group $\operatorname{SO}(2)\times\operatorname{Sol}(2),$ locally isometric to the the three-dimensional subspace $(S_{0},ds_{0}^{2})$ of the Gödel Universe. According to (43) and (10), we have
$W_{1}=[Y,Z]=\sqrt{2}[-f_{1}/2,(f_{0}-f_{2})]=-\sqrt{2}f_{2},\quad\gamma_{1}=-\frac{1}{2}(X,W_{1})=\sqrt{2}.$
So, according to previous calculations,
$R(X,Y)Y=-2X,\quad(R(X,Y)Y,X)=-2.$ (38)
Let us compute now $R(X,Y)Y$ for left-invariant Lorentz metric on $\operatorname{SL}(2,\mathbb{R}),$ defined in [3]. It follows from (36) that
$[X,Y]=\sqrt{2}Z,\quad[X,Z]=\sqrt{2}Y,\quad[Y,Z]=2\sqrt{2}X.$ (39)
Let us apply (34) and (39) in the further computations.
$(\nabla_{X}Y,Z)=-\sqrt{2},\quad(\nabla_{X}Y,X)=(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=\sqrt{2}Z,$
$(\nabla_{Y}Y,X)=(\nabla_{Y}Y,Y)=(\nabla_{Y}Y,Z)=0,\quad\nabla_{Y}Y=0,$
$(\nabla_{Z}Y,X)=-2\sqrt{2},\quad(\nabla_{Z}Y,Y)=(\nabla_{Z}Y,Z)=0,\quad\nabla_{Z}Y=-2\sqrt{2}X,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=0,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=-\sqrt{2}(\nabla_{Y}Z-\nabla_{Z}Y)=-4X.$ (40)
All equalities (37), (38), and (40) are pairwise different.
This implies all statements of Theorem 3. ∎
6. Isometry of non-isomorphic sub-Riemannian Lie groups
Let us change notation $x_{1}\leftrightarrow x_{2},$ $x_{0}\rightarrow x_{3}.$ Then the mapping
$\left(\begin{array}[]{cccc}e^{-x_{2}}&0&0&x_{1}\\ 0&1&0&x_{2}\\ 0&0&1&x_{3}\\ 0&0&0&1\end{array}\right)\rightarrow\left(\begin{array}[]{ccc}e^{-x_{2}}&0&x_{1}\\ 0&1&x_{3}\\ 0&0&1\end{array}\right)\in A^{+}(\mathbb{R})\times(\mathbb{R},+)$ (41)
is an isomorphism of matrix Lie groups with the basis of the Lie algebra
$e_{1}=\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 0&0&0\end{array}\right),e_{2}=\left(\begin{array}[]{ccc}-1&0&0\\ 0&0&0\\ 0&0&0\end{array}\right),e_{3}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&1\\ 0&0&0\end{array}\right),$
such that only $[e_{1},e_{2}]=-[e_{2},e_{1}]=e_{1}$ are unique nonzero Lie brackets.
Analogously to (19) and (20), the mapping
$F\left(\left(\begin{array}[]{ccc}e^{-x_{2}}&0&x_{1}\\ 0&1&x_{3}\\ 0&0&1\end{array}\right)\right)=\left(\left(\begin{array}[]{cc}e^{-x_{2}/2}&x_{1}\\ 0&e^{x_{2}/2}\end{array}\right),\left(\begin{array}[]{cc}\cos x_{3}&\sin x_{3}\\ -\sin x_{3}&\cos x_{3}\end{array}\right)\right)$ (42)
is the universal covering epimorphism of $A^{+}(\mathbb{R})\times(\mathbb{R},+)$ onto $\operatorname{Sol}(2)\times\operatorname{SO}(2).$
The standard left-invariant sub-Riemannian structure on $A^{+}(\mathbb{R})\times(\mathbb{R},+)$ is defined in [4] by the orthonormal frame $\Delta=span\{e_{2},e_{1}+e_{3}\}.$ Then there is unique sub-Riemannian structure on $\operatorname{Sol}(2)\times\operatorname{SO}(2)$ such that $F$ is a local isometry; it is defined by the orthonormal frame $\overline{\Delta}=span\{\overline{e_{2}},\overline{e_{1}}+\overline{e_{3}}\},$ where
$\overline{e_{1}}=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right),\quad\overline{e_{2}}=\left(\begin{array}[]{cc}-1/2&0\\ 0&1/2\end{array}\right),\quad\overline{e_{3}}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right).$ (43)
Now we follow [4]. Let $a=e^{-x_{2}}$ and $b=x_{1}$ in the second matrix of (41).
The subgroup $A^{+}(\mathbb{R})$ is diffeomorphic to the half-plane $\{(a,b)\in\mathbb{R}^{2},a>0\},$ which is desrcibed in the standard polar coordinates as $\{(\rho,\theta)|\rho>0,-\pi/2<\theta<\pi/2\}.$
Theorem 4.
[4]. The diffeomorphism $\Psi:A^{+}(\mathbb{R})\times\operatorname{S^{1}}^{1}\rightarrow\operatorname{SL}(2,\mathbb{R})$ defined by
$\Psi(\rho,\theta,\varphi)=\frac{1}{\sqrt{\rho\cos\theta}}\left(\begin{array}[]{cc}\cos\varphi&\sin\varphi\\ \rho\sin(\theta-\varphi)&\rho\cos(\theta-\varphi)\end{array}\right),$ (44)
where $(\rho,\theta)\in A^{+}(\mathbb{R})$ and $\varphi\in\operatorname{S^{1}}^{1},$ is a global sub-Riemannian isometry.
Remark 1.
Using the above locally isometric covering $F$, we can and will understand $\Psi$ as the global isometry between $\operatorname{Sol}(2)\times\operatorname{SO}(2)$ and $\operatorname{SL}(2,\mathbb{R})$ supplied with sub-Riemannian metrics defined by the same frame $\overline{\Delta}.$
Corollary 4.
$A^{+}(\mathbb{R})\times(\mathbb{R},+)$ with sub-Riemannian metric, defined by the frame $\Delta,$ is isometric to the universal covering $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ of $\operatorname{SL}(2,\mathbb{R})$ with sub-Riemannian metric such that the natural universal covering epimorphism of $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ onto $\operatorname{SL}(2,\mathbb{R})$ with sub-Riemannian metric, defined by the frame $\overline{\Delta},$ is a local isometry.
Proposition 3.
The global isometry $\Psi$ in the sense of Remark 1 is the Iwasawa diffeomorphism of $\operatorname{Sol}(2)\times\operatorname{SO}(2)$ onto $\operatorname{SL}(2,\mathbb{R})$ of the view $(n\overline{a},k)\in NA\times\operatorname{SO}(2)\rightarrow n\overline{a}k\in NAK=\operatorname{SL}(2,\mathbb{R})$, where
$n=\left(\begin{array}[]{cc}1&0\\ b&1\end{array}\right),\overline{a}=\left(\begin{array}[]{cc}a^{-1/2}&0\\ 0&a^{1/2}\end{array}\right),k=\left(\begin{array}[]{cc}\cos\varphi&\sin\varphi\\ -\sin\varphi&\cos\varphi\end{array}\right),a=\rho\cos\theta,b=\rho\sin\theta.$
Proof.
One needs simply to check that $n\overline{a}k$ is equal to the matrix in (44). ∎
Remark 2.
Notice that $n=\exp(t\tilde{e}_{1}),$ $\overline{a}=\exp(s\overline{e_{2}}),$ where $\tilde{e}_{1}=(\overline{e}_{1})^{T},$ T is the sign of transposition, $b=t$, and $a^{1/2}=e^{s/2}$. Also $[\tilde{e}_{1},\overline{e_{2}}]=-\tilde{e}_{1}.$