### v2

# УДК 514.82 + 512.81 + 517.977+517.838+519.46

Abstract.
We discuss models of the Gödel Universe as Lie groups with left-invariant Lorentz metric for two simply connected four-dimensional Lie groups, the Iwasawa decomposition for semisimple Lie groups, and left-invariant Lorentz metric on $\operatorname{SL}(2,\mathbb{R})$, following K.-H. Neeb. Also we show that the isometry between two non-isomorphic sub-Riemannian Lie group, constructed by A. Agrachev and D. Barilari, is induced by some Iwasawa decomposition of $\operatorname{SL}(2,\mathbb{R})$.
2020 Mathematical Subject Classification 83C20, 53C17, 53C50, 49J15, 22E30
Keywords and phrases: Gödel Universe, Iwasawa decomposition, left-invariant Lorentz metric, left-invariant sub-Riemannian metric, Lie algebra, Lie group.
The work was carried out within the framework of the State Contract to the IM SB RAS, project FWNF-2022-0006.
Dedicated to the memory of Vitaly Roman’kov
1. Introduction
Kurt Gödel in paper [1] of 1949 introduced the Lorentz metric (1) of the signature (+,,,)(+,-,-,-)( + , - , - , - ) on the space $\mathbb{R}^{4}.$ The Gödel Universe (space-time) $S$ is a solution of the General Relativity Theory (the Einstein gravitation equations).
In paper [2], the author found timelike and isotropic geodesics of the Gödel Universe, considered as the Lie group $G=(\mathbb{R},+)\times\operatorname{A}^{+}(\mathbb{R})\times(\mathbb{R},+)$ with left invariant Lorentz metric. Here $\operatorname{A}^{+}(\mathbb{R})$ is connected Lie group of affine transformations on $\mathbb{R}.$
The above left-invariant Lorentz metric on $G$ and behaviour of its geodesics are defined essentially by the corresponding induced left-invariant Lorentz metric $ds^{2}_{0}$ on the subgroup $G_{3}=(\mathbb{R},+)\times\operatorname{A}^{+}(\mathbb{R}).$
Professor Karl–Hermann Neeb wrote to the author that it is possible to realize the Gödel Universe otherwise. He sent an electronic version of his joint with Joachim Hilgert book [3], where in section 2.7 ’’Gödel’s cosmological model and universal covering of $\operatorname{SL}(2,\mathbb{R})$’’ is suggested a left-invariant Lorentz metric on $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ which is isometric to $(G_{3},ds^{2}_{0})$ as stated there.
In this connection, it is useful to mention the paper [4] by A. Agrachev and D. Barilari, where the autors obtained a full classification of left-invariant sub-Riemannian metrics on three-dimensional Lie groups and ’’explicitly find a sub-Riemannian isometry between nonisomorphic Lie groups $\operatorname{SL}(2,\mathbb{R})$ and $\operatorname{SO}(2)\times A^{+}(\mathbb{R})$’’ [4].
The existence of such isometry was indicated ealier in [5] by Falbel and Gorodski.
In a message to the author, Professor Neeb explains the mentioned two isometries by a diffeomormism of Lie groups $\operatorname{SL}(2,\mathbb{R})$ and $\operatorname{SO}(2)\times A^{+}(\mathbb{R})$ by means of the Iwasawa decomposition for $\operatorname{SL}(2,\mathbb{R})$; let us cite Theorems 6.5.1 and 9.1.3 from [6].
In p. 10.6.4 (i) from [6] are indicated the following isomorphisms of Lie algebras:
$\mathfrak{sl}(2,\mathbb{R})\cong\mathfrak{su}(1,1)\cong\mathfrak{so}(2,1)\cong\mathfrak{sp}(1,\mathbb{R}).$
Consequently, simply connected Lie groups with these Lie algebras are isomorphic.
In Theorem 3 we prove some properties of special left-invariant Lorentz metrics on Lie groups which support the mentioned statement on the isometry of two left-invariant Lorentz metrics from [3]. Also we show in Proposition 3 that the isometry between two non-isomorphic sub-Riemannian Lie group, constructed by A. Agrachev and D. Barilari, is induced by some Iwasawa decomposition of $\operatorname{SL}(2,\mathbb{R})$.
The author expresses his gratitudes to Professor Neeb for fruitful discussions and anonymous referee for useful remarks.
2. The Gödel Universe as a Lie group with left-invariant Lorentz metric
Gödel introduced in [1] his space-time $S$ as $\mathbb{R}^{4}$ with the linear element
$ds^{2}=a^{2}\left(dx_{0}^{2}+2e^{x_{1}}dx_{0}dx_{2}+\frac{e^{2x_{1}}}{2}dx_{2}^{2}-dx_{1}^{2}-dx_{3}^{2}\right),\quad a>0.Ыд$ (1)
Gödel noticed in [1] that it is possible to rewrite this quadratic form in view of
$ds^{2}=a^{2}\left[\left(dx_{0}+e^{x_{1}}dx_{2}\right)^{2}-dx^{2}_{1}-\frac{e^{2x_{1}}}{2}dx^{2}_{2}-dx^{2}_{3}\right],$ (2)
which shows obvious that its signature is equal everywhere to (+,,,).(+,-,-,-).( + , - , - , - ) .
We shall assume that $a=1.$
Gödel noticed in [1] that on $(S,ds^{2})$ acts simply transitively a four-dimensional isometry Lie group. It is easy to see that such action could be written as
$x_{0}=x_{0}^{\prime}+a,\quad x_{1}=x_{1}^{\prime}+b,\quad x_{2}=x_{2}^{\prime}e^{-b}+c,\quad x_{3}=x_{3}^{\prime}+d$ (3)
with arbitrary $a,b,c,d\in\mathbb{R}.$ This implies that corresponding Lie group $G$ is the simplest simply connected noncommutative four-dimensional Lie group of the view
$G\cong[(\mathbb{R},+)\times G_{2}]\times(\mathbb{R},+),$ (4)
where $G_{2}$ is unique up to isomorphism, necessary isomorphic to $\mathbb{R}^{2}$, two-dimensional noncommutative Lie group. The Lie group $G_{2}$ is isomorphic to the Lie group $A^{+}(\mathbb{R})$ of preserving orientation affine transformations of the real direct line $(\mathbb{R},+).$
[…]
$A=\left(\begin{array}[]{ccccc}e^{-b}&0&0&0&c\\ 0&1&0&0&b\\ 0&0&1&0&a\\ 0&0&0&1&d\\ 0&0&0&0&1\end{array}\right).$ (5)
Under this the equality
$\left(\begin{array}[]{ccccc}e^{-x_{1}}&0&0&0&x_{2}\\ 0&1&0&0&x_{1}\\ 0&0&1&0&x_{0}\\ 0&0&0&1&x_{3}\\ 0&0&0&0&1\end{array}\right)(0,0,0,0,1)^{T}=(x_{2},x_{1},x_{0},x_{3},1)^{T}$ (6)
sets the bijection of the group $G$ onto $\mathbb{R}^{4}$ and the unit of $G$ corresponds to the zero-vector $(0,0,0,0)\in\mathbb{R}^{4}.$ On base of this, (4) and (1), we can identify $(S,ds^{2})$ with the Lie group $G$ equipped with left-invariant Lorentz metric. Let
$e_{0}=\frac{\partial}{\partial x_{0}}(0),e_{1}=\frac{\partial}{\partial x_{1}}(0),e_{2}=\frac{\partial}{\partial x_{2}}(0),e_{3}=\frac{\partial}{\partial x_{3}}(0))$
be the basis of the Lie algebra $\mathfrak{g}$ of the Lie group $G$ at the unit of $G$, corresponding to coordinates $(x_{0},x_{1},x_{2},x_{3})$. Then, according to what has been said and (1), the components of the linear element $ds^{2}$ with respect to this basis are equal to
$g_{00}=1,g_{02}=g_{20}=1,g_{22}=\frac{1}{2},g_{11}=-1,g_{33}=-1,g_{ij}=g_{ji}=0,i\neq j,j\neq 2.$ (7)
According to (6), the Lie subgroup $G_{3}:=(\mathbb{R},+)\times G_{2}$ can be identified with the matrix Lie group
$\left(\begin{array}[]{cccc}e^{-x_{1}}&0&0&x_{2}\\ 0&1&0&x_{1}\\ 0&0&1&x_{0}\\ 0&0&0&1\end{array}\right),\quad(x_{0},x_{1},x_{2})\in\mathbb{R}^{3}.$ (8)
It is obvious that $(S,ds^{2})=(S_{0},ds_{0}^{2})\times(S_{1},ds_{1}^{2}),$ where $S_{0}=\mathbb{R}^{3},$ $S_{1}=\mathbb{R},$
$ds_{0}^{2}=dx_{0}^{2}+2e^{x_{1}}dx_{0}dx_{2}+\frac{e^{2x_{1}}}{2}dx_{2}^{2}-dx_{1}^{2},\quad ds_{1}^{2}=-dx_{3}^{2}.$ (9)
Also it is clear that we can consider $(S_{0},ds_{0}^{2})$ as the matrix Lie group (8) with left-invariant Lorentz metric, which according to (7) has components
$g_{00}=1,g_{02}=g_{20}=1,g_{22}=\frac{1}{2},g_{11}=-1,g_{ij}=g_{ji}=0,i\neq j,j\neq 2;$ (10)
with respect to the basis $e_{0},e_{1},e_{2}$ of the Lie algebra $\mathfrak{g}_{3}$ of the matrix Lie group (8).
In consequence of (6), for the Lie algebra $\mathfrak{g}_{3}$ of the Lie group $G_{3},$
$e_{0}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\end{array}\right),\quad e_{1}=\left(\begin{array}[]{cccc}-1&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\end{array}\right),\quad e_{2}=\left(\begin{array}[]{cccc}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right).$ (11)
Then in the Lie algebra $\mathfrak{g}_{3}$,
$[e_{1},e_{2}]=e_{1}e_{2}-e_{2}e_{1}=-e_{2},\quad[e_{0},e_{1}]=[e_{0},e_{2}]=0.$ (12)
3. The Iwasawa decompositions of Lie algebras and Lie groups
Let $\mathfrak{g}$ be a semisimple real Lie algebra, $\sigma$ be some Cartan involution of $\mathfrak{g},$ and $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be the corresponding Cartan decomposition ($\mathfrak{k}$ is the Lie subalgebra of $\mathfrak{g}$, consisting of fixed points relative to $\sigma$). Let us denote by $\mathfrak{a}$ a maximal commutative subspace in $\mathfrak{p}.$ Then there is the following Iwasawa decomposition of Lie algebra $\mathfrak{g}$.
Theorem 1.
(4.7.2) in [7]. Let $\mathfrak{g}$ be a semisimple real Lie algebra.
Then there exists a direct sum of vector subspaces in $\mathfrak{g}$
$\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{a}\oplus\mathfrak{n},$ (13)
where $\mathfrak{n}$ is a nilpotent subalgebra in $\mathfrak{g}$ such that the endomorphism $\operatorname{ad}X$ is nilpotent for every $X\in\mathfrak{n},$ and $\mathfrak{a}\oplus\mathfrak{n}$ is a solvable subalgebra in $\mathfrak{g}.$
As an example, the authors of [7] give the decomposition (13) for $\mathfrak{g}=\mathfrak{sl}(n,\mathbb{R}).$ In this case $\mathfrak{k}$ is the Lie subalgebra of skew-symmetric matrices, $\mathfrak{a}$ is the Lie subalgebra of diagonal matrices with zero trace, and $\mathfrak{n}$ is the Lie subalgebra of strictly upper triangular matrices. In particular, for $\mathfrak{g}=\mathfrak{sl}(2,\mathbb{R})$ we have
$\mathfrak{k}=\left\{\left(\begin{array}[]{cc}0&t\\ -t&0\end{array}\right)\right\},\quad\mathfrak{a}=\left\{\left(\begin{array}[]{cc}t&0\\ 0&-t\end{array}\right)\right\},\quad\mathfrak{n}=\left\{\left(\begin{array}[]{cc}0&t\\ 0&0\end{array}\right)\right\},\quad t\in\mathbb{R},$ (14)
with natural basis
$f_{0}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),\quad f_{1}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),\quad f_{2}=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right)$ (15)
and Lie brackets for this basis
$[f_{0},f_{1}]=2f_{0}-4f_{2},\quad[f_{0},f_{2}]=f_{1},\quad[f_{1},f_{2}]=2f_{2}.$ (16)
Let $K=\exp(\mathfrak{k}),$ $A=\exp(\mathfrak{a}),$ $N=\exp(\mathfrak{n})$ be Lie subgroups of the semisimple Lie group $G,$ corresponding to the decomposition (13).
Theorem 2.
(Theorem 9.1.3 in [6]) Let $G$ be a connected semisimple real Lie group. Then $G=KAN$ and the mapping
$(k,a,n)\rightarrow kan$ (17)
is the diffeomorphism of manifold $K\times A\times N$ onto the Lie group $G.$
Corollary 1.
The Lie group $G$ is diffeomorphic to Lie groups $K\times AN$ and $K\times A\times N.$
Theorem 1 and Theorem 6.1.1 from [6] imply the following
Proposition 1.
The sets $K,$ $A,$ $N,$ and $AN$ are connected closed Lie subgroups of the Lie group $G,$ where $\operatorname{Ad}_{G}(K)$ is compact, $A$ is commutative, $N$ is nilpotent, and $AN$ is solvable. The subgroup $K$ contains the center $Z$ of the Lie group $G.$ In addition, $K$ is compact if and only if the center $Z$ of $G$ is finite; in this case $K$ is a maximal compact subgroup of the Lie group $G.$
The statements above imply the following
Corollary 2.
If $G=\operatorname{SL}(n,\mathbb{R}),$ then $K=\operatorname{SO}(n),$ $A$ is the group of all real diagonal $(n\times n)$-matrices with unit determinant, $N$ could be considered as the group of all real upper triangular $(n\times n)$-matrices with units on the main diagonal, and $\operatorname{Sol}(n):=AN$ as the group of all real upper triangular $(n\times n)$-matrices with unit determinant.
Corollary 3.
The Lie group $\operatorname{SL}(n,\mathbb{R})$ is diffeomorphic to Lie groups $\operatorname{SO}(n)\times AN$ and $\operatorname{SO}(n)\times A\times N.$ As a consequence, $\operatorname{SL}(2,\mathbb{R})$ is diffeomorphic to the commutative Lie group $\operatorname{SO}(2)\times A\times N.$
4. $(S_{0},ds^{2}_{0})$ as (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric
This is a preparatory section.
Proposition 2.
There exist an isomorphism of the Lie group $G_{3}$ onto the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ and corresponding realization of $(S_{0},ds^{2}_{0})$ as the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric.
Proof.
Comparing (12) and (16), we see that the linear map $\varphi:\mathfrak{a}\oplus\mathfrak{n}\rightarrow\mathfrak{g}_{2}$ such that
$\varphi\left(\frac{-f_{1}}{2}\right)=e_{1},\quad\varphi(f_{2})=e_{2}$ (18)
is an isomorphism of Lie algebras. Let $\operatorname{Sol}(2)$ be the Lie group with the Lie algebra $\mathfrak{a}\oplus\mathfrak{n}$ for (14). Then (18) defines isomorphism of Lie groups $\psi:\operatorname{Sol}(2)\rightarrow G_{2}:$
$\psi\left(\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)\right)=\left(\begin{array}[]{cccc}e^{-s}&0&0&e^{-s}r\\ 0&1&0&s\\ 0&0&1&0\\ 0&0&0&1\end{array}\right),$ (19)
$\psi\left(\left(\begin{array}[]{cc}1&r\\ 0&1\end{array}\right)\left(\begin{array}[]{cc}e^{-s/2}&0\\ 0&e^{s/2}\end{array}\right)\right)=\left(\begin{array}[]{cccc}e^{-s}&0&0&r\\ 0&1&0&s\\ 0&0&1&0\\ 0&0&0&1\end{array}\right).$ (20)
We can consider $(\mathbb{R},+)$ as a Lie algebra and as a Lie group. Then the mappings
$t\in(\mathbb{R},+)\rightarrow\left(\begin{array}[]{cc}0&t\\ -t&0\end{array}\right),\quad t\in(\mathbb{R},+)\rightarrow\left(\begin{array}[]{cc}\cos t&\sin t\\ -\sin t&\cos t\end{array}\right)$ (21)
are correspondingly the isomorphism of Lie algebras and respective universal covering epimorpism of Lie groups. Then there exists unique isomorphism $\psi$ of the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ onto the Lie group $G_{3},$ with properties (19), (20), and $\psi(t)=t$ for $t\in(\mathbb{R},+).$ It follows from previous considerations that we shall realize $(S_{0},ds^{2}_{0})$ as the Lie group (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$ with left-invariant Lorentz metric if components of this metric in the basis $\{f_{0},-f_{1}/2,f_{2}\}$ of its Lie algebra will be as in (10).
The corresponding orthonormal basis is
$X=f_{0},\quad Y^{\prime}=-f_{1}/2,\quad Z^{\prime}=\sqrt{2}(f_{0}-f_{2}),$ (22)
which we can change by
$X=f_{0},\quad Y=f_{1}/2,\quad Z=\sqrt{2}(f_{2}-f_{0}).$ (23)
Then
$[Y,Z]=[Y^{\prime},Z^{\prime}]=\sqrt{2}f_{2}=Z+\sqrt{2}X.$ (24)
Remark 1.
The basis (23) has a form
$X=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),Y=\frac{1}{2}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),Z=\sqrt{2}\left(\begin{array}[]{cc}0&0\\ 1&0\end{array}\right).$ (25)
The group $\operatorname{Sol}(2)$ is isomorphic to the Lie group of real lower triangular $(2\times 2)-$matrices with unit determinant.
5. Left-invariant Lorentz metrics on $\operatorname{SO}(2)\times\operatorname{Sol}(2)$ and $\operatorname{SL}(2,\mathbb{R})$
In [3], the authors consider the Lie group $\operatorname{SL}(2,\mathbb{R})$ with left-invariant Lorentz metric and orthonormal basis with the signature (,+,+)(-,+,+)( - , + , + ) of the form
$X=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),Y=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right),Z=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\in\mathfrak{sl}(2,\mathbb{R}).$ (26)
We shall consider this basis as orthonormal with the signature (+,,).(+,-,-).( + , - , - ) .
Theorem 3.
1) For the Lorentz metric on $\operatorname{SL}(2,\mathbb{R})$ from [3] with orthonormal basis (26) and for corresponding by the Iwasawa diffeomorphism basis on $\operatorname{SO}(2)\times\operatorname{Sol}(2)$, the curvatures $k(X,Y)=k(X,Z)=k(Y,Z)=-2,$ while for the orthonormal basis (23) on (,+)×$(\mathbb{R},+)\times\operatorname{Sol}(2)$, $k(X,Y)=k(X,Z)=k(Y,Z)=-\frac{1}{2}$.
2) One-parameter subgroups in $\operatorname{SL}(2,\mathbb{R}),$ defined by $X,$ $Y,$ $Z,$ are geodesics.
Proof.
For any (pseudo-)Riemannian manifold $M$ with (pseudo-)metric tensor (,),($\cdot$,$\cdot$),( ⋅ , ⋅ ) , the Levi-Civita connection $\nabla$, and smooth vector fields $X,$ $Y,$ $Z$ is valid the following equation (3.5.(7)) from [8]:
$(\nabla_{X}Y,Z)=\frac{1}{2}[X(Y,Z)+Y(Z,X)-Z(X,Y)+(Z,[X,Y])+(Y,[Z,X])-(X,[Y,Z])].$ (27)
As a consequence, if $(M,(\cdot,\cdot))$ is a Lie group $G$ with left-invariant (pseudo-)metric (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) of the signature (+,,)(+,-,-)( + , - , - ) and $X,$ $Y,$ $Z$ are left-invariant, then
$(\nabla_{X}Y,Z)=\frac{1}{2}[(Z,[X,Y])+(Y,[Z,X])-(X,[Y,Z])].$ (28)
It follows from (26) that
$[X,Y]=-\sqrt{2}Z,\quad[X,Z]=\sqrt{2}Y,\quad[Y,Z]=2\sqrt{2}X.$ (29)
Let us apply (28) and (29) in the further computations.
$(\nabla_{X}Y,Z)=0,\quad(\nabla_{X}Y,X)=(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=0,$
$(\nabla_{Y}Y,X)=(\nabla_{Y}Y,Y)=(\nabla_{Y}Y,Z)=0,\quad\nabla_{Y}Y=0,$
$(\nabla_{Z}Y,X)=-\sqrt{2},\quad(\nabla_{Z}Y,Y)=(\nabla_{Z}Y,Z)=0,\quad\nabla_{Z}Y=-\sqrt{2}X,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=\sqrt{2}X,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=\sqrt{2}\nabla_{Z}Y=-2X.$ (30)
Then $k(X,Y)=(R(X,Y)Y,X)=-2.$ Analogously, we obtain
$\nabla_{X}Z=0,\quad\nabla_{X}X=0,\quad\nabla_{Z}Z=0,\quad k(X,Z)=(R(X,Z)Z,X)=-2,$
$R(Y,Z)Z=\nabla_{Y}\nabla_{Z}Z-\nabla_{Z}\nabla_{Y}Z-\nabla_{[Y,Z]}Z=-\sqrt{2}\nabla_{Z}X=-\sqrt{2}(\nabla_{X}Z+[Z,X])=2Y,$
$k(Y,Z)=(R(Y,Z)Z,Y)=-2.$
2) $[Y,Z]=2Z+2\sqrt{2}X$ in $\mathfrak{so}(2)\oplus\mathfrak{sol}(2).$
We look for $Z$ as $\alpha f_{2}+\beta X$, see (15). It is easy to see that $\alpha=2,$ $\beta=-\sqrt{2},$
$[Y,Z]=[Y,2f_{2}-\sqrt{2}X]=2[f_{1},f_{2}]=4f_{2}=2(Z+\sqrt{2}X)=2Z+2\sqrt{2}X.$
$(\nabla_{X}Y,Z)=-\sqrt{2},\quad(\nabla_{X}Y,X)=(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=\sqrt{2}Z,$
$(\nabla_{Y}Y,X)=(\nabla_{Y}Y,Y)=(\nabla_{Y}Y,Z)=0,\quad\nabla_{Y}Y=0,$
$(\nabla_{Z}Y,X)=-\sqrt{2},\quad(\nabla_{Z}Y,Y)=0,\quad(\nabla_{Z}Y,Z)=2,\quad\nabla_{Z}Y=-\sqrt{2}X-2Z,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=\sqrt{2}X,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=-\sqrt{2}\nabla_{Y}Z=-2X.$
Then $k(X,Y)=(R(X,Y)Y,X)=-2.$ Analogously, we obtain
$\nabla_{X}Z=-\sqrt{2}Y,\nabla_{Z}X=\nabla_{X}Z=-\sqrt{2}Y,\quad\nabla_{Z}Z=2Y,\quad k(X,Z)=(R(X,Z)Z,X)=-2,$
$R(Y,Z)Z=\nabla_{Y}\nabla_{Z}Z-\nabla_{Z}\nabla_{Y}Z-\nabla_{[Y,Z]}Z=$
$\nabla_{Y}(2Y)-\nabla_{Z}(\sqrt{2}X)-\nabla_{(2Z+2\sqrt{2}X)}Z=$
$2Y-2\nabla_{Z}Z-2\sqrt{2}\nabla_{X}Z=2Y-4Y+4Y=2Y,$
$k(Y,Z)=(R(Y,Z)Z,Y)=-2.$
Let us compute analogous expressions for $(S_{0},ds_{0}^{2}),$ using results of Section 4.
$(\nabla_{X}Y,Z)=-\frac{\sqrt{2}}{2},\quad(\nabla_{X}Y,X)=(\nabla_{X}Y,Y)=0,\quad\nabla_{X}Y=\frac{\sqrt{2}}{2}Z,$
$(\nabla_{Y}Y,X)=(\nabla_{Y}Y,Y)=(\nabla_{Y}Y,Z)=0,\quad\nabla_{Y}Y=0,$
$(\nabla_{Z}Y,X)=-\frac{\sqrt{2}}{2},\quad(\nabla_{Z}Y,Y)=0,\quad(\nabla_{Z}Y,Z)=1,\quad\nabla_{Z}Y=-\frac{\sqrt{2}}{2}X-Z,$
$\nabla_{Y}Z=\nabla_{Z}Y+[Y,Z]=\frac{\sqrt{2}}{2}X,$
$R(X,Y)Y=\nabla_{X}\nabla_{Y}Y-\nabla_{Y}\nabla_{X}Y-\nabla_{[X,Y]}Y=-\frac{\sqrt{2}}{2}\nabla_{Y}Z=-\frac{1}{2}X.$
Then $k(X,Y)=(R(X,Y)Y,X)=-\frac{1}{2}.$ Analogously, we obtain
$\nabla_{X}Z=-\frac{\sqrt{2}}{2}Y,\nabla_{Z}X=\nabla_{X}Z=-\frac{\sqrt{2}}{2}Y,\quad\nabla_{Z}Z=Y,\quad k(X,Z)=(R(X,Z)Z,X)=-\frac{1}{2},$
$R(Y,Z)Z=\nabla_{Y}\nabla_{Z}Z-\nabla_{Z}\nabla_{Y}Z-\nabla_{[Y,Z]}Z=$
$\nabla_{Y}(Y)-\nabla_{Z}\left(\frac{\sqrt{2}}{2}X\right)-\nabla_{(Z+\sqrt{2}X)}Z=$
$\frac{1}{2}Y-\nabla_{Z}Z-\sqrt{2}\nabla_{X}Z=\frac{1}{2}Y-Y+Y=\frac{1}{2}Y,$
$k(Y,Z)=(R(Y,Z)Z,Y)=-\frac{1}{2}.$
Obtained equalities imply all statements of Theorem 3. ∎
Remark 2.
It is similar, that the left-invariant Lorentz metric on $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ from [3] is isometric to Gödel metric $(S_{0},ds^{2}_{0})$ induced by Gödel metric (1) with $a=\frac{1}{\sqrt{2}}.$
6. Isometry of non-isomorphic sub-Riemannian Lie groups
Let us change notation $x_{1}\leftrightarrow x_{2},$ $x_{0}\rightarrow x_{3}.$ Then the mapping
$\left(\begin{array}[]{cccc}e^{-x_{2}}&0&0&x_{1}\\ 0&1&0&x_{2}\\ 0&0&1&x_{3}\\ 0&0&0&1\end{array}\right)\rightarrow\left(\begin{array}[]{ccc}e^{-x_{2}}&0&x_{1}\\ 0&1&x_{3}\\ 0&0&1\end{array}\right)\in A^{+}(\mathbb{R})\times(\mathbb{R},+)$ (31)
is an isomorphism of matrix Lie groups with the basis of the Lie algebra
$e_{1}=\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 0&0&0\end{array}\right),e_{2}=\left(\begin{array}[]{ccc}-1&0&0\\ 0&0&0\\ 0&0&0\end{array}\right),e_{3}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&1\\ 0&0&0\end{array}\right),$
such that only $[e_{1},e_{2}]=-[e_{2},e_{1}]=e_{1}$ are unique nonzero Lie brackets.
Analogously to (19) and (20), the mapping
$F\left(\left(\begin{array}[]{ccc}e^{-x_{2}}&0&x_{1}\\ 0&1&x_{3}\\ 0&0&1\end{array}\right)\right)=\left(\left(\begin{array}[]{cc}e^{-x_{2}/2}&x_{1}\\ 0&e^{x_{2}/2}\end{array}\right),\left(\begin{array}[]{cc}\cos x_{3}&\sin x_{3}\\ -\sin x_{3}&\cos x_{3}\end{array}\right)\right)$ (32)
is the universal covering epimorphism of $A^{+}(\mathbb{R})\times(\mathbb{R},+)$ onto $\operatorname{Sol}(2)\times\operatorname{SO}(2).$
The standard left-invariant sub-Riemannian structure on $A^{+}(\mathbb{R})\times(\mathbb{R},+)$ is defined in [4] by the orthonormal frame $\Delta=span\{e_{2},e_{1}+e_{3}\}.$ Then there is unique sub-Riemannian structure on $\operatorname{Sol}(2)\times\operatorname{SO}(2)$ such that $F$ is a local isometry; it is defined by the orthonormal frame $\overline{\Delta}=span\{\overline{e_{2}},\overline{e_{1}}+\overline{e_{3}}\},$ where
$\overline{e_{1}}=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right),\quad\overline{e_{2}}=\left(\begin{array}[]{cc}-1/2&0\\ 0&1/2\end{array}\right),\quad\overline{e_{3}}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right).$ (33)
Now we follow [4]. Let $a=e^{-x_{2}}$ and $b=x_{1}$ in the second matrix of (31).
The subgroup $A^{+}(\mathbb{R})$ is diffeomorphic to the half-plane $\{(a,b)\in\mathbb{R}^{2},a>0\},$ which is desrcibed in the standard polar coordinates as $\{(\rho,\theta)|\rho>0,-\pi/2<\theta<\pi/2\}.$
Theorem 4.
[4]. The diffeomorphism $\Psi:A^{+}(\mathbb{R})\times\operatorname{S^{1}}^{1}\rightarrow\operatorname{SL}(2,\mathbb{R})$ defined by
$\Psi(\rho,\theta,\varphi)=\frac{1}{\sqrt{\rho\cos\theta}}\left(\begin{array}[]{cc}\cos\varphi&\sin\varphi\\ \rho\sin(\theta-\varphi)&\rho\cos(\theta-\varphi)\end{array}\right),$ (34)
where $(\rho,\theta)\in A^{+}(\mathbb{R})$ and $\varphi\in\operatorname{S^{1}}^{1},$ is a global sub-Riemannian isometry.
Remark 3.
Using the above locally isometric covering $F$, we can and will understand $\Psi$ as the global isometry between $\operatorname{Sol}(2)\times\operatorname{SO}(2)$ and $\operatorname{SL}(2,\mathbb{R})$ supplied with sub-Riemannian metrics defined by the same frame $\overline{\Delta}.$
Corollary 4.
$A^{+}(\mathbb{R})\times(\mathbb{R},+)$ with sub-Riemannian metric, defined by the frame $\Delta,$ is isometric to the universal covering $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ of $\operatorname{SL}(2,\mathbb{R})$ with sub-Riemannian metric such that the natural universal covering epimorphism of $\widetilde{\operatorname{SL}(2,\mathbb{R})}$ onto $\operatorname{SL}(2,\mathbb{R})$ with sub-Riemannian metric, defined by the frame $\overline{\Delta},$ is a local isometry.
Proposition 3.
The global isometry $\Psi$ in the sense of Remark 3 is the Iwasawa diffeomorphism of $\operatorname{Sol}(2)\times\operatorname{SO}(2)$ onto $\operatorname{SL}(2,\mathbb{R})$ of the view $(n\overline{a},k)\in NA\times\operatorname{SO}(2)\rightarrow n\overline{a}k\in NAK=\operatorname{SL}(2,\mathbb{R})$, where
$n=\left(\begin{array}[]{cc}1&0\\ b&1\end{array}\right),\overline{a}=\left(\begin{array}[]{cc}a^{-1/2}&0\\ 0&a^{1/2}\end{array}\right),k=\left(\begin{array}[]{cc}\cos\varphi&\sin\varphi\\ -\sin\varphi&\cos\varphi\end{array}\right),a=\rho\cos\theta,b=\rho\sin\theta.$
Proof.
One needs simply to check that $n\overline{a}k$ is equal to the matrix in (34). ∎
Remark 4.
Notice that $n=\exp(t\tilde{e}_{1}),$ $\overline{a}=\exp(s\overline{e_{2}}),$ where $\tilde{e}_{1}=(\overline{e}_{1})^{T},$ T is the sign of transposition, $b=t$, and $a^{1/2}=e^{s/2}$. Also $[\tilde{e}_{1},\overline{e_{2}}]=-\tilde{e}_{1}.$